# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

### v1

@@ 00 $\ell$
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ $1$
$\lambda_{\ell-2}$ $1$
@@ 00 $\ell(\ell+1)/2$
We shall present the computations for $i=2$ and we may remark that the case $i=3$ follows very similarly by resorting to the analogous arguments. So, we put $\lambda=\lambda_{\ell-2}$ and $\mu=\lambda_{\ell-1}$. Since $\mu=\lambda-(\alpha_{1}+\alpha_{2}+\cdots+\alpha_{\ell-2})$ we have to consider in the Freudenthal formula the positive roots of the form
@@
Indeed, the formula for $m_{\lambda}(\mu)$ involves the multiplicities of $\mu+(\alpha_{i}+\cdots+\alpha_{j})$ for any $21\leq i\leq j\leq\ell-$ 2. We then observe
@@
Now, we claim that the positive roots $\alpha_{1}+\alpha_{2}+\cdots+\alpha_{k}$, with $21\leq k\leq\ell-$ 2, are the only ones that contribute in the computation of $m_{\lambda}(\mu)$. Indeed, we can show that $m_{\lambda}(\mu+(\alpha_{i}+\cdots+\alpha_{j}))=0$, for any $22\leq i\leq j\leq\ell-$ 2. To see that, we directly verify that $m_{\lambda}(\lambda-\alpha_{k})=0$ for any $31\leq k\leq\ell-$ 3 which implies inductively that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1}))=0$, i.e, we prove the case $j=\ell-2$. Now, by assuming now that $j<\ell-2$ we have $\mu+(\alpha_{i}+\cdots+\alpha_{j})=\lambda-(\alpha_{1}+\cdots+\alpha_{i-1})+(\alpha_{j+1}+\cdots+\alpha_{\ell-2}).$
Then we can note that the kind of positive roots that contribute in the Fruedental formula to compute the foregoing multiplicity are of the form $\alpha_{i^{\prime}}+\cdots+\alpha_{j^{\prime}}$ and $\alpha_{j}$, with $i\leq i^{\prime}\leq j^{\prime}\leq j$, so we may deduce that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1}+\alpha_{j}))=0$ for any $i+1\leq j\leq\ell-2$. By combining these multiplicities we can conclude that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1})-(\alpha_{j+1}+\cdots+\alpha_{t}))=0$, for $j+1\leq t\leq\ell-2$. In summary, we have proved that
@@
and we write $\mu_{k}=\lambda-(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{\ell-2})$.
By arguing as above and thanks to (5.2) only the weights of the form $\mu_{k}+(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{j})$ will contribute in the Fruedenthal formula of $m_{\lambda}(\lambda-(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{\ell-2}))$. So, we observe
@@
[…]
@@
At this point, we need to remark that $m_{\lambda}(\mu+2(\alpha_{1}+\cdots+\alpha_{k}))(\mu+2(\alpha_{1}+\cdots+\alpha_{k}),(\alpha_{1}+\cdots+\alpha_{k}))$ also contributes in the Freudenthal formula of $m_{\lambda}(\mu)$, when $31\leq k\leq\ell-$ 3. According to the terms associated to the above positive roots in the foregoing formula, we first compute that
@@
Supposing that $\alpha_{k}=\alpha_{1}+\cdots+\alpha_{k}$, where $11\leq k\leq\ell-$ 1, we directly verify that $(\mu+\alpha_{k},\alpha_{k})=1$. Now we shall prove that $m_{\lambda}(\mu+\alpha_{k})=1$. Let us start by noticing that the multiplicities for $k=\ell-i,\;i=1,2$ were computed in the above case, thus we have to compute $m_{\lambda}(\mu+\alpha_{k})$ for $k\leq\ell-3$. For the general case, the Freudenthal formula involves computing many terms of the form $m_{\lambda}(\mu_{k}+\alpha))(\mu_{k}+\alpha,\alpha)$ for appropriate positive roots $\alpha$ we start by setting $\mu_{k}=\mu+\alpha_{k}$. Each of them implies in its turn to compute many other multiplicities. Nevertheless, we introduce a case sufficiently illustrative of these computations, say $m_{\lambda}(\mu_{\ell-3})$. Firstly, we see that
@@
Regarding the term $m_{\lambda}(\mu+2\alpha_{k})(\mu+2\alpha_{k},\alpha_{k})$, for $31\leq k\leq\ell-$ 3, we can proceed similarly as above to deduce each multiplicity $m_{\lambda}(\mu+2\alpha_{k})$. At this point, routine computations allow to calculate the remaining components to obtain the multiplicities.
Finally, suppose that $\alpha_{k,j}=2(\alpha_{1}+\cdots+\alpha_{k})+(\alpha_{k+1}+\cdots+\alpha_{j})$, where $31\leq k\leq\ell-$ 3 and $\ell-2\leq j\leq\ell-1$. We compute $(\mu+\alpha_{k,j},\alpha_{k,j})$ as follows:
@@
    $=(\lambda_{\ell-2},\alpha_{\ell-2})-2(\alpha_{k+1},\alpha_{k})-(\alpha_{k+1},\alpha_{k+1})-(\alpha_{\ell-2},\alpha_{\ell-1})$
    $=2.$
@@
    $=(\lambda_{\ell-2},\alpha_{\ell-2})-2(\alpha_{k+1},\alpha_{k})-(\alpha_{k+1},\alpha_{k+1})-(\alpha_{\ell-2},\alpha_{\ell-1})$
    $=2,$
@@
On the other hand, we may resort as above to determine the only positive roots $\alpha$ involving in the formula of $m_{\lambda}(\mu+\alpha_{k,j})$ we recursively deduce that $m_{\lambda}(\lambda_{\ell-2}-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))=1$ as in the previous case. Since $31\leq k\leq\ell-$ 3 and $\ell-2\leq j\leq\ell-1$ it follows that the number of positive roots $\alpha_{k,j}$ is
@@
In summary, the computations from the cases above provide respectively the following summands to the formula of $m_{\lambda}(\mu)$
12,(1)11,(3)12,(2)12, $1\cdot 2,\;(\ell-1)\cdot 1\cdot 1,\;(\ell-3)\cdot 1\cdot 2,\;(\ell-2)\cdot 1\cdot 2,\;1\cdot 2(\ell-2)(\ell-3)$
Then we can compute
$m_{\lambda}(\mu)$ $=\frac{2[2\cdot 1+(\ell-1)\cdot 1+2(\ell-2)+2(\ell-3)+2(\ell-2)(\ell-3)]}{2(2\ell-3)}$
$=\frac{2(\ell-1)(2\ell-3)}{2(2\ell-3)}$
$=\ell-1.$
@@
This writing allows to apply the same strategy as above to determine all terms involved in the formula of $m_{\lambda}(0)$ and thereby to conclude the desired multiplicity.
@@
For all $\ell\geq 12$, the multiplicities of all dominant weights $\mu\preceq\lambda_{\ell-1}+\lambda_{\ell}$ are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ multiplicity of $\mu$
$\lambda=\lambda_{\ell-1}+\lambda_{\ell}$ $\lambda_{\ell-1}+\lambda_{\ell}$ $1$
$\lambda_{\ell-2}$ $2$
$2\lambda_{\ell}$ $1$
$\lambda_{\ell-1}$ $2$
$\lambda_{\ell}$ $12\ell-$ 1
00 $12\ell-$ 1
@@
We shall start by computing the multiplicity of $\lambda_{\ell-2}=\lambda-(\alpha_{\ell-1}+\alpha_{\ell})$. To do this, we can verify as above that
$m_{\lambda}(\lambda-\alpha_{\ell})=2\cdot\frac{m_{\lambda}((\lambda-\alpha_{\ell})+\alpha_{\ell})((\lambda-\alpha_{\ell})+\alpha_{\ell},\alpha_{\ell})}{(\lambda,\lambda)-(\mu_{\ell},\mu_{\ell})+2(\alpha_{\ell},\delta)}=2\cdot\frac{1\cdot 1}{5-5+2\cdot 1}=1.$
and similarly we check that $m_{\lambda}(\lambda-\alpha_{\ell-1})=1$. Thus, we may deduce that the multiplicity $m_{\lambda}(\lambda_{\ell-2})=2$ by computing as follows
$m_{\lambda}(\lambda_{\ell-2})$ $=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell-1})(\mu_{\ell-1}+\alpha_{\ell-1},\alpha_{\ell-1})+m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell})(\mu_{\ell-1}+\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda,\alpha_{\ell-1}+\alpha_{\ell})}{(\lambda,\lambda)-(\lambda_{\ell-2},\lambda_{\ell-2})+2(\alpha_{\ell-1}+\alpha_{\ell},\delta)}$
$=2\cdot\frac{1\cdot 2+1\cdot 2+1\cdot 2}{5-3+4}=2.$
In order to compute $m_{\lambda}(2\lambda_{\ell})$ we see that
$\mu:=2\lambda_{\ell}=\lambda-\sum_{k=1}^{\ell-1}\alpha_{k}.$
A direct computation shows
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-1}\alpha_{k},\delta)=5-4+2(1/2+(\ell-2))=2(\ell-1).$
By arguing as in proof of Lemma 5.3, we may infer that the only required multiplicities to compute $m_{\lambda}(2\lambda_{\ell})$ are $m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))$, for $21\leq k\leq\ell-$ 2, and $m_{\lambda}(\lambda)$ itself. Inductively, as Lemma 5.4, we may check that $m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))=1$ for any $21\leq k\leq\ell-$ 2. In addition, we can also verify in this case that
$(\mu+(\alpha_{1}+\cdots+\alpha_{k}),\alpha_{1}+\cdots+\alpha_{k})$ $=(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-1}),\alpha_{1}+\cdots+\alpha_{k})=-(\alpha_{k+1},\alpha_{k})=1.$
for $k\leq\ell-2$. Thus, we deduce
$m_{\lambda}(\mu)$ $=2\cdot\frac{\displaystyle{\sum_{k=1}^{\ell-1}m_{\lambda}(\mu+(\alpha_{1}+\cdots+\alpha_{k}))(\mu+(\alpha_{1}+\cdots+\alpha_{k}),\alpha_{1}+\cdots+\alpha_{k})}}{(\lambda,\lambda)-(\mu,\mu)-2(\sum_{k=1}^{\ell-1}\alpha_{k},\delta)}$
$=2\cdot\frac{\displaystyle{\sum_{k=1}^{\ell-1}m_{\lambda}(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}),\alpha_{1}+\cdots+\alpha_{k})}}{2(\ell-1)}$
$=2\cdot\frac{\ell-1}{2(\ell-1)}=1.$
We now start the computations of $m_{\lambda}(\lambda_{\ell-1})$. We first note that
$\mu:=\lambda_{\ell-1}=\lambda-\sum_{k=1}^{\ell}\alpha_{k}.$
We have
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell}\alpha_{k},\delta)=5-2+2(1/2+(\ell-1))=2(\ell+1).$
In addition we already know that $m_{\lambda}(\lambda-(\alpha_{\ell-1}+\alpha_{\ell}))=2$ and $m_{\lambda}(\lambda-\alpha_{\ell})=1$ by calculations above. Thus, we see that $m_{\lambda}(\lambda)(\lambda,\alpha_{1}+\cdots+\alpha_{\ell})=1\cdot 2$. Indeed, we may argue as in Lemma 5.3 to deduce
$m_{\lambda}(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell}))(\lambda-(\alpha_{k+1}+\cdots+\alpha_{\ell}),\alpha_{1}+\cdots+\alpha_{k})=2\cdot 1,$
for any $21\leq k\leq\ell-$ 2, and $m_{\lambda}(\lambda)(\lambda,\alpha_{1}+\cdots+\alpha_{\ell})=1\cdot 2$. In the same way, we can verify
$m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{\ell-1}))(\lambda-(\alpha_{1}+\cdots+\alpha_{\ell-1}),\alpha_{\ell})=1\cdot 2.$
Therefore, by noting that the foregoing terms are the only nonzero summands in the Fruedenthal formula as in Lemma 5.4 we conclude that $m_{\lambda}(\lambda_{\ell-1})=2$.
While the multiplicity $m_{\lambda}(\lambda_{\ell})$ may be computed by means of the arguments used in Lemma 5.4 observing that
$\mu:=\lambda_{\ell}=\lambda-\sum_{k=1}^{\ell-1}\alpha_{k}-\sum_{k=1}^{\ell}\alpha_{k}=\lambda-2\sum_{k=1}^{\ell-1}\alpha_{k}-\alpha_{\ell}$
and
$(\lambda,\lambda)-(\mu,\mu)+2(2\sum_{k=1}^{\ell-1}\alpha_{k}+\alpha_{\ell},\delta)=5-1+4(\ell-1)=4\ell.$
Idem for $m_{\lambda}(0)$ taking in mind
$\mu:=0=\lambda-\sum_{k=1}^{\ell-1}\alpha_{k}-2\sum_{k=1}^{\ell}\alpha_{k}=3(2\ell-1).$
and
$(\lambda,\lambda)-(\mu,\mu)+2(\sum_{k=1}^{\ell-1}\alpha_{k}+2\sum_{k=1}^{\ell}\alpha_{k},\delta)=5-0+3(2\ell-1))=2(3\ell+1).$
Thus we can prove.
@@
Let $G$ be as above of type $B_{\ell}$ with $\ell\geq 12$. Then, all weights in the following table are admissible.
$\lambda$ dimension of $V(\lambda)$
00 $1$
$\lambda_{\ell}$ $12\ell+$ 1
$2\lambda_{\ell}$ $\ell(2\ell+3)-1$
$3\lambda_{\ell}$ $\ell(2\ell+1)(2\ell+5)/3$
$4\lambda_{\ell}$ $\ell(\ell+1)(2\ell+1)(2\ell+7)/6$
$\lambda_{\ell-1}$ $2\ell^{2}+\ell$
$\lambda_{\ell-2}$ $(2\ell-1)2\ell(2\ell+1)/6$
$\lambda_{\ell-3}$ $(2\ell-2)(2\ell-1)2\ell(2\ell+1)/24$
$\lambda_{\ell}+\lambda_{\ell-1}$ $(2\ell+3)(2\ell+1)(2\ell-1)/3$
In addition, when $12\leq\ell\leq 15$ the table has to include the weight $\lambda_{1}$ whose associated irreducible representation $V(\lambda_{1})$ has dimension $2^{\ell}$.

### v3

@@ 0 $\ell$
$\lambda_{\ell-3}$ $\lambda_{\ell-3}$ 11
$\lambda_{\ell-2}$ 11
@@ 0 $\ell(\ell+1)/2$
We shall present the computations for $i=2$ and we may remark that the case $i=3$ follows very similarly by resorting to the analogous arguments. So, we put $\lambda=\lambda_{\ell-2}$ and $\mu=\lambda_{\ell-1}$. Since $\mu=\lambda-(\alpha_{1}+\alpha_{2}+\cdots+\alpha_{\ell-2})$ we have to consider in the Freudenthal formula (2.2), the positive roots of the form
@@
Indeed, the formula for $m_{\lambda}(\mu)$ involves the multiplicities of $\mu+(\alpha_{i}+\cdots+\alpha_{j})$ for any $1\leq i\leq j\leq\ell-2$. We then observe
@@
Now, we claim that the positive roots $\alpha_{1}+\alpha_{2}+\cdots+\alpha_{k}$, with $1\leq k\leq\ell-2$, are the only ones that contribute in the computation of $m_{\lambda}(\mu)$. Indeed, we can show that $m_{\lambda}(\mu+(\alpha_{i}+\cdots+\alpha_{j}))=0$, for any $2\leq i\leq j\leq\ell-2$. To see that, we directly verify that $m_{\lambda}(\lambda-\alpha_{k})=0$ for any $1\leq k\leq\ell-3$ which implies inductively that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1}))=0$, i.e, we prove the case $j=\ell-2$. Now, by assuming now that $j<\ell-2$ we have $\mu+(\alpha_{i}+\cdots+\alpha_{j})=\lambda-(\alpha_{1}+\cdots+\alpha_{i-1})+(\alpha_{j+1}+\cdots+\alpha_{\ell-2}).$ Then we can note that the kind of positive roots that contribute in the Fruedental formula to compute the foregoing multiplicity are of the form $\alpha_{i^{\prime}}+\cdots+\alpha_{j^{\prime}}$ and $\alpha_{j}$, with $i\leq i^{\prime}\leq j^{\prime}\leq j$, so we may deduce that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1}+\alpha_{j}))=0$ for any $i+1\leq j\leq\ell-2$. By combining these multiplicities we can conclude that $m_{\lambda}(\lambda-(\alpha_{1}+\cdots+\alpha_{i-1})-(\alpha_{j+1}+\cdots+\alpha_{t}))=0$, for $j+1\leq t\leq\ell-2$. In summary, we have proved that
@@
and we write $\mu_{k}=\lambda-(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{\ell-2})$. By arguing as above and thanks to (4.2) only the weights of the form $\mu_{k}+(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{j})$ will contribute in the Fruedenthal formula of $m_{\lambda}(\lambda-(\alpha_{k-1}+\alpha_{k}+\cdots+\alpha_{\ell-2}))$. So, we observe
@@
$\alpha$ $=\alpha_{1}+\cdots+\alpha_{k},\;\quad\>\;1\leq k\leq\ell-1,$
$\alpha$ $=\alpha_{k}+\cdots+\alpha_{\ell-1},\;\>\;2\leq k\leq\ell-1,$
@@
At this point, we need to remark that $m_{\lambda}(\mu+2(\alpha_{1}+\cdots+\alpha_{k}))(\mu+2(\alpha_{1}+\cdots+\alpha_{k}),(\alpha_{1}+\cdots+\alpha_{k}))$ also contributes in the Freudenthal formula of $m_{\lambda}(\mu)$, when $1\leq k\leq\ell-3$. According to the terms associated to the above positive roots in the foregoing formula, we first compute that
@@
Supposing that $\alpha_{k}=\alpha_{1}+\cdots+\alpha_{k}$, where $1\leq k\leq\ell-1$, we directly verify that $(\mu+\alpha_{k},\alpha_{k})=1$. Now we shall prove that $m_{\lambda}(\mu+\alpha_{k})=1$. Let us start by noticing that the multiplicities for $k=\ell-i,\;i=1,2$ were computed in the above case, thus we have to compute $m_{\lambda}(\mu+\alpha_{k})$ for $k\leq\ell-3$. For the general case, the Freudenthal formula involves computing many terms of the form $m_{\lambda}(\mu_{k}+\alpha))(\mu_{k}+\alpha,\alpha)$ for appropriate positive roots $\alpha$ we start by setting $\mu_{k}=\mu+\alpha_{k}$. Each of them implies in its turn to compute many other multiplicities. Nevertheless, we introduce a case sufficiently illustrative of these computations, say $m_{\lambda}(\mu_{\ell-3})$. Firstly, we see that
@@
Regarding the term $m_{\lambda}(\mu+2\alpha_{k})(\mu+2\alpha_{k},\alpha_{k})$, for $1\leq k\leq\ell-3$, we can proceed similarly as above to deduce each multiplicity $m_{\lambda}(\mu+2\alpha_{k})$. At this point, routine computations allow to calculate the remaining components to obtain the multiplicities.
Finally, suppose that $\alpha_{k,j}=2(\alpha_{1}+\cdots+\alpha_{k})+(\alpha_{k+1}+\cdots+\alpha_{j})$, where $1\leq k\leq\ell-3$ and $\ell-2\leq j\leq\ell-1$. We compute $(\mu+\alpha_{k,j},\alpha_{k,j})$ as follows:
@@
    $=(\lambda_{\ell-2},\alpha_{\ell-2})-2(\alpha_{k+1},\alpha_{k})-(\alpha_{k+1},\alpha_{k+1})-(\alpha_{\ell-2},\alpha_{\ell-1})=2.$
@@
    $=(\lambda_{\ell-2},\alpha_{\ell-2})-2(\alpha_{k+1},\alpha_{k})-(\alpha_{k+1},\alpha_{k+1})-(\alpha_{\ell-2},\alpha_{\ell-1})=2,$
@@
On the other hand, we may resort as above to determine the only positive roots $\alpha$ involving in the formula of $m_{\lambda}(\mu+\alpha_{k,j})$. Then, we recursively deduce that $m_{\lambda}(\lambda_{\ell-2}-(\alpha_{k+1}+\cdots+\alpha_{\ell-2}))=1$ as in the previous case. Since $1\leq k\leq\ell-3$ and $\ell-2\leq j\leq\ell-1$, it follows that the number of positive roots $\alpha_{k,j}$ is
@@
In summary, the computations from the cases above provide respectively the following summands to the formula of $m_{\lambda}(\mu)$:
$1\cdot 2,\quad(\ell-1)\cdot 1\cdot 1,\quad(\ell-3)\cdot 1\cdot 2,\quad(\ell-2)\cdot 1\cdot 2,\quad 1\cdot 2(\ell-2)(\ell-3)$
Then, we have that
$m_{\lambda}(\mu)=\frac{2[2\cdot 1+(\ell-1)\cdot 1+2(\ell-2)+2(\ell-3)+2(\ell-2)(\ell-3)]}{2(2\ell-3)}=\ell-1.$
@@
This writing allows to apply the same strategy as above to determine all terms involved in the formula of $m_{\lambda}(0)$ and thereby to conclude the required multiplicity. Then, the desired bound follows from (2.3) and (2.4)).
@@
For all $\ell\geq 12$ and $p\neq 2$, a bound for the multiplicities of all dominant weights $\mu\preceq i\lambda_{\ell}$, $i=3,4$, are given in the following table:
highest weight dominant weight $\mu\preceq\lambda$ a bound of $m_{L(\lambda)}(\mu)$
$3\lambda_{\ell}$ $3\lambda_{\ell}$ 11
$\lambda_{\ell-1}+\lambda_{\ell}$ 11
$\lambda_{\ell-2}$ 11
$2\lambda_{\ell}$ 11
$\lambda_{\ell-1}$ 11
$\lambda_{\ell}$ $\ell$
0 $\ell$
$4\lambda_{\ell}$ $4\lambda_{\ell}$ 11
$\lambda_{\ell-1}+2\lambda_{\ell}$ 11
$2\lambda_{\ell-1}$ 11
$\lambda_{\ell-2}+\lambda_{\ell}$ 11
$\lambda_{\ell-3}$ 11
$3\lambda_{\ell}$ 11
$\lambda_{\ell-1}+\lambda_{\ell}$ 11
$\lambda_{\ell-2}$ 11
$2\lambda_{\ell}$ $\ell$
$\lambda_{\ell-1}$ $\ell$
$\lambda_{\ell}$ $\ell$
0 $\ell(\ell-1)/2$
@@
Let $\lambda=3\lambda_{\ell}$ and $\mu:=\lambda_{\ell-1}+\lambda_{\ell}=\lambda-\alpha_{\ell}$. Then, we have that $(\lambda,\lambda)=9$ and $(\mu,\mu)=5$. In addition as $(\alpha_{\ell},\delta)=1$ and $(\lambda,\alpha_{\ell})=3(\lambda_{\ell},\alpha_{\ell})=1$, we have
Next, compute the multiplicity of $\lambda_{\ell-2}=\lambda-(\alpha_{\ell-1}+2\alpha_{\ell})$, we observe that the positive roots involved in the Fruedenthal formula are $\alpha_{\ell-1}$, $\alpha_{\ell}$, $2\alpha_{\ell}$ and $\alpha_{\ell-1}+\alpha_{\ell}$. Notice, at this point, that $\lambda$ is not included in the formula since $\mu+\alpha_{\ell-1}+2\alpha_{\ell}$ is not of the form $\mu+i\alpha$ for some positive root $\alpha$ by appealing to [6, Plate II].
Set $\mu:=\lambda-(\alpha_{\ell-1}+\alpha_{\ell})$.
  1. (1)
    $m_{\lambda}(\mu_{\ell-1})=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell})(\mu_{\ell-1}+\alpha_{\ell},\alpha_{\ell})}{(\lambda,\lambda)-(\mu_{\ell-1},\mu_{\ell-1})+2(\alpha_{\ell},\delta)}=1.$
  2. (2)
$m_{\lambda}(\mu)$ $=2\cdot\frac{m_{\lambda}(\mu_{\ell-1}+\alpha_{\ell-1})(\mu_{\ell-1}+\alpha_{\ell-1},\alpha_{\ell-1})+m_{\lambda}(\mu_{\ell}+\alpha_{\ell})(\mu_{\ell}+\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda_{\ell},\alpha_{\ell-1}+\alpha_{\ell})}{(\lambda,\lambda)-(\lambda_{\ell-2},\lambda_{\ell-2})+2(\alpha_{\ell-1}+\alpha_{\ell},\delta)}=1.$
$m_{\lambda}(\lambda-2\alpha_{\ell})$ $=2\cdot\frac{m_{\lambda}(\lambda-\alpha_{\ell})(\lambda-\alpha_{\ell},\alpha_{\ell})+m_{\lambda}(\lambda)(\lambda,\alpha_{\ell})}{(\lambda,\lambda)-(\mu,\mu)+2(2\alpha_{\ell},\delta)}=1.$
where
$(*)$ $=(\mu+\alpha_{\ell},\alpha_{\ell})m_{\lambda}(\mu+\alpha_{\ell})+(\mu+\alpha_{\ell-1})m_{\lambda}(\mu+\alpha_{\ell-1},\alpha_{\ell-1})$
$=(\mu+\alpha_{\ell},\alpha_{\ell})\cdot 1+(\mu+\alpha_{\ell-1},\alpha_{\ell-1})\cdot 1+(\mu+2\alpha_{\ell},\alpha_{\ell})\cdot 0+(\mu+\alpha_{\ell-1}+\alpha_{\ell},\alpha_{\ell-1}+\alpha_{\ell})\cdot 1$
$=2+2+0+2=6.$
Computing as above we see that
$m_{\lambda}(\lambda_{\ell-2})=\frac{2[1\cdot 2+1\cdot 2+0\cdot 4+1\cdot 2]}{12}=1.$
The remaining multiplicities $m_{\lambda}(\mu)$, for $\mu\preceq 2\lambda_{\ell}$, can be computed as in [24].
Finally, the case $\lambda=4\lambda_{\ell}$ is longer but completely similar than the one above by calculating the multiplicity of $\lambda_{\ell-1}+2\lambda_{\ell}$ as first step. ∎
@@
Let $G$ be as above of type $B_{\ell}$ with $\ell\geq 12$ and assume $p\neq 2$. The following weights
$3\lambda_{\ell}\;4\lambda_{\ell},\;\lambda_{\ell-2},\;\lambda_{\ell-3},\;\lambda_{\ell-1}+\lambda_{\ell}$
are admissible weights. Additionally, $\lambda_{1}$ is also an admissible weight when $12\leq\ell\leq 16$.